What is the maximum amount you would pay for an asset that generates an income of $250,000 at the end of each of five years, if the opportunity cost of using funds is 8 percent?
Do not round intermediate calculations. Round only your final calculations to the nearest penny( two decimal places). Use a negative sign (-) where appropriate.
Answer #1

ANSWER:

I = 8%

N = 5

INCOME = $250,000

Pv = income(p/a,i,n) = 250,000(p/a,8%,5) = 250,000 * 3.993 = $998,250

so the maximum amount that i would pay for the asset is $998,250